Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by heshamelaziry » Mon Nov 16, 2009 9:58 pm
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

OA [spoiler]16/21[/spoiler]
Join the discussion
Source: — Problem Solving |

by Abdulla » Mon Nov 16, 2009 11:12 pm
heshamelaziry wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

OA [spoiler]16/21[/spoiler]
I am bad with Probabilities but I will try it..

Prob = # desire outcomes / Total outcomes

Let's start with the easy one. We want to choose 2 out of 7.

Total outcomes = 7! / 2!5! = 21 ( 7 x 6 x 5 x 4 x 3 x 2 x 1 / 5 x 4 x 3 x 2 x 1 x 2 x 1 )

# of Desired outcomes. We've got two scenarios.
1) Choosing two people from the first group.
4! / 2!2! = 6 ..... ( remember that we've selected 2 people each of which has one sibling in the room, so we should subtract 2)
6-2 = 4 .......(1)

2) Choosing one from the first group times another from the second group.
4! /1!3! * 3! /2!1! = 4*3 = 12 ...........(2)

4+12 / 21 = 16/21 #
Abdulla
Join the discussion

by heshamelaziry » Mon Nov 16, 2009 11:34 pm
HOhw you say you are bad with probability ? you got the right answer. However, I am still lost.
Join the discussion

by Abdulla » Mon Nov 16, 2009 11:49 pm
heshamelaziry wrote:HOhw you say you are bad with probability ? you got the right answer. However, I am still lost.
I spent 15 minutes to figure it out. :D
Abdulla
Join the discussion

by heshamelaziry » Mon Nov 16, 2009 11:55 pm
Abdulla wrote:
heshamelaziry wrote:HOhw you say you are bad with probability ? you got the right answer. However, I am still lost.
I spent 15 minutes to figure it out. :D

Still very well done. I still can't rationalize it :x . BTW, did u get my pm ?
Join the discussion

by getso » Tue Nov 17, 2009 4:19 am
Hi ,

Could you please explain the logic...
Join the discussion

by viju9162 » Tue Nov 17, 2009 4:32 am
Hi Abdulla, heshamelaziry,

I didnt understand how you arrived at the answer. Can you explain how can we solve with this formula:

P (A) = 1 - P`(A).

Regards,
Viju
"Native of" is used for a individual while "Native to" is used for a large group
Join the discussion

by mridul_dave » Tue Nov 17, 2009 6:35 am
heshamelaziry wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

OA [spoiler]16/21[/spoiler]
Lets say letters represents a person and a '--' represents the sibling relation.

A -- B
B -- C
these are the 4 people who have exactly one sibling.


E -- F
\ /
G

There are the three people with exactly two siblings each. We have total 7 people.

We just made three sets of people.

Lets get the prob. that two are siblings. For that either you pick from the first set, or second set or any two from the third set.

2C2 + 2 C2 + 3C2 = 1 + 1 + 3 = 5

Total ways is 7C2 = 21. Prob that two are siblings is 5/21.
So prob they are not is 1 - 5/21 = 16/21.
Join the discussion

by getso » Tue Nov 17, 2009 6:51 am
Thanks for the great explaination ..

:)
Join the discussion

by Abdulla » Tue Nov 17, 2009 11:29 pm
mridul_dave wrote:
heshamelaziry wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

OA [spoiler]16/21[/spoiler]
Lets say letters represents a person and a '--' represents the sibling relation.

A -- B
B -- C
these are the 4 people who have exactly one sibling.


E -- F
\ /
G

There are the three people with exactly two siblings each. We have total 7 people.

We just made three sets of people.

Lets get the prob. that two are siblings. For that either you pick from the first set, or second set or any two from the third set.

2C2 + 2 C2 + 3C2 = 1 + 1 + 3 = 5

Total ways is 7C2 = 21. Prob that two are siblings is 5/21.
So prob they are not is 1 - 5/21 = 16/21.
WOW, Easy approach..
Thanks hesham
Abdulla
Join the discussion